# SeaS: Few-shot Industrial Anomaly Image Generation with Separation and Sharing Fine-tuning

### v1

@@ 4.3 Ablation Study
Anomaly image generation. We train additional models to assess the effect of each component: (a) the model without mixing the different types of anomaly images in the same product; (b) the model without NA loss; (c) the model without the second term of DA loss in Eq. 3 (short for ST in Tab. 5); (d) our complete model. We use these models to generate 1,000 anomaly image-mask pairs per anomaly type and train BiSeNet V2 for anomaly segmentation. In Tab. 5, the results show that omitting any component leads to a decrease in the fidelity and diversity of the generated images, as well as a decrease in the segmentation results. These validate the effectiveness of the components we proposed. More ablation studies on SeaS are shown in appendix A.8.
Refined Mask Prediction branch. To verify the validity of the components in the RMP branch, we conduct ablation studies on MRM, the progressive manner to refine coarse feature (short for PM in Tab. 5) and coarse mask supervision (short for CMS in Tab. 5). 1) the model without any components, which means we do not use MRM to fuse the high-resolution features in RMP, but directly obtain the mask from the coarse features $\hat{F}\in\mathbb{R}^{64\times 64\times 192}$ through convolution and bilinear interpolation upsampling; 2) the model with MRM; 3) the model utilizing the MRM in a progressive manner to refine coarse features; 4) our complete model. We report the BiSeNet V2 results in Tab. 5, which demonstrates that each component in the RMP is indispensable for downstream anomaly segmentation. More ablation studies about RMP are in appendix A.8.
@@
(a) w/o Mixed 1.79 0.32 95.82 66.07 64.50 53.11
(b) w/o NA 1.67 0.31 96.20 66.03 64.09 53.97
(c) w/o ST 1.86 0.33 96.44 67.73 65.23 54.99
(d) All (Ours) 1.88 0.34 97.21 69.21 66.37 55.28
@@ 5 Conclusion
In this paper, we propose a novel few-shot industrial anomaly image generation method named SeaS. We explore an implicit characteristic that the anomalies exhibit randomness in shape and appearance, while the products maintain global consistency with minor variations in local details. We design a Separation and Sharing Fine-tuning strategy for industrial anomaly image generation, and a Refined Mask Prediction branch to obtain a fine-grained mask. Our method surpasses existing methods on both AIG and downstream anomaly segmentation tasks.
6 Acknowledgments
@@ A.1 Overview
  • Anomaly image and mask generation results, downstream segmentation results on MVTec 3D AD dataset (RGB images) (Sec. A.4).
  • More quantitative comparison with unsupervised anomaly detection methods (Sec. A.5).
@@
  • More details of downstream segmentation model implementation and usage (Sec. A.7).
  • Qualitative comparison results of segmentation models trained on image-mask pairs generated by different anomaly generation methods (Sec. A.9).
  • Qualitative comparison results of different segmentation models trained on image-mask pairs generated by SeaS (Sec. A.10).
@@ A.2 Analysis on decoupled Anomaly alignment loss and multiple tokens
Here we give a more detailed analysis of the learning process of the DA loss. According to Eq. 3, intuitively, the DA loss may pull the anomaly tokens similar to each other. However, the U-Net in Stable Diffusion uses multi-head attention, which ensures different anomaly tokens cover different attributes of the anomalies. In Eq. 3, the cross-attention map is the multiply of the feature map of U-Net and the anomaly tokens. In the implementation of multi-head attention, both the learnable embedding of the anomaly token and the U-Net feature are decomposed into several groups along the channel dimension. E.g., the conditioning vector ${e}_{a}\in\mathbb{R}^{1\times C_{1}}$, which is corresponding to anomaly token, is divided into ${\{{e}_{a,i}\in\mathbb{R}^{1\times\frac{C_{1}}{q}}|i\in[1,q]\}}$, and the image feature ${v}\in\mathbb{R}^{r\times r\times C_{2}}$ is divided into ${\{{v}_{i}\in\mathbb{R}^{1\times\frac{C_{2}}{q}}|i\in[1,q]\}}$, where $q$ is the number of heads in the multi-head attention. Then the corresponding groups are multiplied, and the outputs of all the heads are averaged. The attention map $A$ of ${e}_{a}$ is calculated by:
@@
Therefore, in the defect region, the DA loss only ensures the average of each head tends to $1$, but does not require the anomaly tokens to be the same with each other. In addition, each ${e}_{a}$ is different from each other, and is combined by ${e}_{a,i}$. The update direction of each ${e}_{a,i}$ is related to ${v}_{i}$ and covers some features of the defect, it encompasses the attributes of anomalies from various perspectives, thereby providing diversified information.
@@ A.3 More implementation details
More training details. For the Unbalanced Abnormal Text Prompt, we set the number $N$ of multiple <$\texttt{df}_{n}$> to $4$ and the number $N^{\prime}$ of <ob> to 1, these parameters are fixed across all product classes. For example, for the normal token <ob>, given the lookup $\mathcal{U}\in\mathbb{R}^{b\times 768}$, where $b$ is the number of text embeddings stored by the pre-trained text encoder, we use a placeholder string "ob1" as the input. Firstly, "ob1" is converted to a token ID $s_{\texttt{ob1}}\in\mathbb{R}^{1\times 1}$in the tokenizer. Secondly, $s_{\texttt{ob1}}\in\mathbb{R}^{1\times 1}$is converted to a one-hot vector $\mathcal{S}_{\texttt{ob1}}\in\mathbb{R}^{1\times(b+1)}$. Thirdly, one learnable new embedding $g\in\mathbb{R}^{1\times 768}$ corresponding to $s_{\texttt{ob1}}$ is inserted to the lookup $\mathcal{U}$, resulting in $\mathcal{U}^{{}^{\prime}}\in\mathbb{R}^{(b+1)\times 768}$. Here, $g\in\mathbb{R}^{1\times 768}$ is the learnable embedding of <ob>. These embeddings and U-Net are learnable during the fine-tuning process.
@@
Training image generation model. For each product, we perform $800\times G$ steps for fine-tuning, where $G$ represents the number of anomaly categories of the product. The batch size of training image generation model is set to 4. During each step of our fine-tuning process, we sample 2 images from the abnormal training set $X_{\text{df}}$, and 2 images from the normal training set $X_{\text{ob}}$. We utilize the AdamW (Loshchilov & Hutter, 2018) optimizer with a learning rate of U-Net is $4\times 10^{-6}$. The learning rate of the text embedding is $4\times 10^{-5}$.
@@
Training Refined Mask Prediction branch. We design a cascaded Refined Mask Prediction (RMP) branch, which is grafted onto the U-Net trained according to SeaS. For each product, we perform $800\times G$ steps for the RMP model, where $G$ represents the number of anomaly types for the product. The batch size of training the RMP branch is set to 4. During each step of our fine-tuning process, we sample 2 images with their corresponding masks from the abnormal training set $X_{\text{df}}$, and 2 images from the normal training set $X_{\text{ob}}$. The masks used to suppress noise in normal images has each pixel value set to 0. The learning rate of the RMP model is $5\times 10^{-4}$.
@@
Metrics. For anomaly image generation, we report 2 metrics: the Inception Score (IS) and Intra-cluster pairwise LPIPS Distance (IC-LPIPS). The Inception Score (IS), proposed in (Barratt & Sharma, 2018), serves as an independent metric to evaluate the fidelity and diversity of generated images, by measuring the mutual information between input samples and their predicted classes. The IC-LPIPS (Ojha et al., 2021) is used to evaluate the diversity of generated images, which quantifies the perceptual similarity between image patches in the same cluster. For pixel-level anomaly segmentation and image-level anomaly detection, we report 3 metrics: Area Under Receiver Operator Characteristic curve (AUROC), Average Precision (AP), and $F_{1}$-score at the optimal threshold ($F_{1}$-max). All of these metric are calculated using the scikit-learn library. In addition, we calculate the Intersection over Union (IoU) to more accurately evaluate the anomaly segmentation result.
@@
Resource requirement and time consumption. We conduct our training on NVIDIA Tesla A100 40G GPU. Specifically, we use a single A100 to train a generation model sequentially for each product category, with each training process occupying about 20G of GPU memory. Since each anomaly type requires isolate training, the training time depends on the total amount of anomaly types across all products. For example, the product metal_nut contains 4 anomaly types, and each needs around 35 minutes. The generation model for metal_nut spends 2 hours and 20 minutes on training in total. For the RMP branch, each anomaly type needs around 25 minutes. Hence, it takes 1 hour and 40 minutes to train metal_nut. More details are given in Tab. 6, where $K$ is the total number of anomaly types across all products. The comparison on time consumption is shown in Tab. 7. For the MVTecAD datasets with 73 anomaly types, our training takes 73 hours, which is shorter than the 249 hours required by AnomalyDiffusion and the 414 hours required by DFMGAN. In terms of inference time, SeaS costs 720 ms per image, which is shorter than the 3830 ms per image required by the Diffusion-based method AnomalyDiffusion. The inference time of the GAN-based method DFMGAN is 48ms per image.
Table 6: Computational resource and training time.
Stage Time (minutes per product) GPU(MB) Overall Time
Generation Model 35 × $K$ 20242 42 hours and 35 minutes
RMP branch 25 × $K$ 23280 30 hours and 25 minutes
Table 7: Comparison on time consumption.
Model Overall Training Time(hours) Inference Time (ms)
DFMGAN (Duan et al., 2023) 414 48
AnomalyDiffusion (Hu et al., 2024) 249 3830
Ours 73 720
A.4 Additional dataset results
We perform experimental evaluations on the RGB images of the MVTec 3D AD Dataset (Bergmann. et al., 2022), which includes 10 product categories, each with up to 4 different anomalies. It encompasses several common challenges, such as variations in lighting conditions and product poses, which are crucial for validating the robustness of image generation methods. The experimental settings are the same as those in Sec. 4.1 and Sec. A.3.

### v2

@@ 4.3 Ablation Study
We train additional models to assess the effect of each component: (a) the model with predefined typical text prompt with fixed generic semantic words (short for with TP in Tab. 5); (b) the model without mixing different types of anomaly images in the same product; (c) the model without NA loss; (d) the model without the second term of DA loss in Eq. 3 (short for ST in Tab. 5); (e) our complete model. We use these models to generate 1,000 anomaly image-mask pairs per anomaly type and train BiSeNet V2 for supervised anomaly segmentation. In Tab. 5, the results show that omitting any component leads to a decrease in fidelity and diversity of the generated images, as well as in the segmentation results. These validate the effectiveness of the components we proposed. More ablation studies on SeaS are shown in appendix A.5.
@@
(a) with TP 1.72 0.33 94.72 57.16 55.67 50.46
(b) w/o Mixed 1.79 0.32 95.82 66.07 64.50 53.11
(c) w/o NA 1.67 0.31 96.20 66.03 64.09 53.97
(d) w/o ST 1.86 0.33 96.44 67.73 65.23 54.99
(e) All (Ours) 1.88 0.34 97.21 69.21 66.37 55.28
Refined Mask Prediction branch. To verify the validity of the components in the RMP branch, we conduct ablation studies on MRM, the progressive manner to refine coarse feature (short for PM in Tab. 6) and coarse mask supervision (short for CMS in Tab. 6). 1) the model only with CMS, which means we do not use MRM to fuse the high-resolution features in RMP, but directly obtain the mask from the coarse features $\hat{F}\in\mathbb{R}^{64\times 64\times 192}$ through convolution and bilinear interpolation upsampling; 2) the model with MRM; 3) the model utilizing three MRMs in a progressive manner to refine coarse features; 4) our complete model. We report the BiSeNet V2 results in Tab. 6, which demonstrates that each component in the RMP is indispensable for downstream supervised anomaly segmentation. More ablation studies about RMP are in appendix A.5.
@@ 5 Conclusion
In this paper, we propose a unified generation method named SeaS. We explore an implicit characteristic that anomalies exhibit high variability, while normal products maintain global consistency. We design a Separation and Sharing Fine-tuning strategy to model different variations of normal products and anomalies, enabling the Refined Mask Prediction branch to predict accurate masks with discriminative features. Our method greatly improves synthesis-based and supervised AD methods, and empowers supervised segmentation models.
6 Acknowledgement
@@ A.1 Overview
  • More details of downstream supervised segmentation model implementation and usage (Sec. A.4).
@@
  • Qualitative comparison results of supervised segmentation models trained on image-mask pairs generated by different anomaly generation methods (Sec. A.7).
  • Qualitative comparison results of different supervised segmentation models trained on image-mask pairs generated by SeaS (Sec. A.8).
  • Comparison with the Textual Inversion (Sec. A.9).
  • More experiments on lighting conditions (Sec. A.10).
  • More results on generation of small defects (Sec. A.11).
  • More analysis on generation of unseen anomaly types (Sec. A.12).
  • More experiments on comparison with DRAEM (Sec. A.13.)
@@ A.2 Analysis on decoupled Anomaly alignment loss and multiple tokens
Here we give a more detailed analysis of the learning process of the DA loss. According to Eq. 3, intuitively, the DA loss may pull the anomaly tokens similar to each other. However, the U-Net in Stable Diffusion uses multi-head attention, which ensures that different anomaly tokens cover different attributes of the anomalies. In Eq. 3, the cross-attention map is the product of the feature map of U-Net and the anomaly tokens. In the implementation of multi-head attention, both the learnable embedding of the anomaly token and the U-Net feature are decomposed into several groups along the channel dimension. E.g., the conditioning vector ${e}_{a}\in\mathbb{R}^{1\times C_{1}}$, which is corresponding to anomaly token, is divided into ${\{{e}_{a,i}\in\mathbb{R}^{1\times\frac{C_{1}}{q}}|i\in[1,q]\}}$, and the image feature ${v}\in\mathbb{R}^{r\times r\times C_{2}}$ is divided into ${\{{v}_{i}\in\mathbb{R}^{1\times\frac{C_{2}}{q}}|i\in[1,q]\}}$, where $q$ is the number of heads in the multi-head attention. Then the corresponding groups are multiplied, and the outputs of all the heads are averaged. The attention map $A$ of ${e}_{a}$ is calculated by:
@@
Therefore, in the defect region, the DA loss only ensures the average of each head tends to $1$, but does not require the anomaly tokens to be the same as each other. In addition, each ${e}_{a}$ is different from each other, and is combined by ${e}_{a,i}$. The update direction of each ${e}_{a,i}$ is related to ${v}_{i}$ and covers some features of the defect, it encompasses the attributes of anomalies from various perspectives, thereby providing diversified information.
We provide more examples in Fig. 8, where new anomalies are generated that significantly differ from the training samples in terms of color and shape. For example, we showcase bottle_contamination, hazelnut_print, and tile_gray_stroke with a novel shape, wood_color and metal_nut_scratch with a novel color, and pill_crack with a new shape, featuring multiple cracks where the training samples only exhibit a single crack. These examples demonstrate the model’s ability to create unseen anomalies based on recombining the decoupled attributes.
Figure 8: Visualization of the generation results for unseen anomalies on MVTec AD.
@@ A.3 More implementation details
More training details. For the Unbalanced Abnormal Text Prompt, we set the number $N$ of multiple <$\texttt{df}_{n}$> to $4$ and the number $N^{\prime}$ of <ob> to 1, these parameters are fixed across all product classes. For a particular type of anomaly, we use the Unbalanced Abnormal (UA) Text Prompt $\mathcal{P}_{n}$ with different sets of anomaly tokens as the condition to generate the specified type of anomaly.
$\mathcal{P}_{n}$ = a <ob> with <$\texttt{df}_{4\times n-3}$>,
<$\texttt{df}_{4\times n-2}$>,<$\texttt{df}_{4\times n-1}$>, <$\texttt{df}_{4\times n}$>
@@
where $n$ represents the index of the anomaly types in the product. To generate normal images, we use the embedding the embedding $\textbf{e}_{\text{ob}}$ corresponding to the normal tokens of $\mathcal{P}$, i.e., “a <ob>”, to guide the U-Net in predicting noise. For example, for the normal token <ob>, given the lookup $\mathcal{U}\in\mathbb{R}^{b\times 768}$, where $b$ is the number of text embeddings stored by the pre-trained text encoder, we use a placeholder string "ob1" as the input. Firstly, "ob1" is converted to a token ID $s_{\texttt{ob1}}\in\mathbb{R}^{1\times 1}$in the tokenizer. Secondly, $s_{\texttt{ob1}}\in\mathbb{R}^{1\times 1}$is converted to a one-hot vector $\mathcal{S}_{\texttt{ob1}}\in\mathbb{R}^{1\times(b+1)}$. Thirdly, one learnable new embedding $g\in\mathbb{R}^{1\times 768}$ corresponding to $s_{\texttt{ob1}}$ is inserted to the lookup $\mathcal{U}$, resulting in $\mathcal{U}^{{}^{\prime}}\in\mathbb{R}^{(b+1)\times 768}$. Here, $g\in\mathbb{R}^{1\times 768}$ is the learnable embedding of <ob>. These embeddings and U-Net are learnable during the fine-tuning process.
@@
Training image generation model. For each product, we perform $800\times G$ steps for fine-tuning, where $G$ represents the number of anomaly categories of the product. The batch size of the training image generation model is set to 4. During each step of our fine-tuning process, we sample 2 images from the abnormal training set $X_{\text{df}}$, and 2 images from the normal training set $X_{\text{ob}}$. We utilize the AdamW [25] optimizer with a learning rate of U-Net is $4\times 10^{-6}$. The learning rate of the text embedding is $4\times 10^{-5}$.
@@
Training Refined Mask Prediction branch. We design a cascaded Refined Mask Prediction (RMP) branch, which is grafted onto the U-Net trained according to SeaS. For each product, we perform $800\times G$ steps for the RMP model, where $G$ represents the number of anomaly types for the product. The batch size of training the RMP branch is set to 4. During each step of our fine-tuning process, we sample 2 images with their corresponding masks from the abnormal training set $X_{\text{df}}$, and 2 images from the normal training set $X_{\text{ob}}$. The masks used to suppress noise in normal images have each pixel value set to 0. The learning rate of the RMP model is $5\times 10^{-4}$.
More inference details. For all experiments, we use $t$ = 1500 to perform diffusion forward on normal images to get the initial noise. We employ $T$ = 25 steps for sampling.
@@
Metrics. For anomaly image generation, we report 4 metrics: the Inception Score (IS) and Intra-cluster pairwise LPIPS Distance (IC-LPIPS) to evaluate the anomaly images, KID [5] to assess the authenticity of normal images, and IC-LPIPS calculated only on anomaly regions (short for IC-LPIPS(a)), to evaluate diversity. The Inception Score (IS), proposed in [2], serves as an independent metric to evaluate the fidelity and diversity of generated images, by measuring the mutual information between input samples and their predicted classes. The IC-LPIPS [29] is used to evaluate the diversity of generated images, which quantifies the perceptual similarity between image patches in the same cluster. For pixel-level anomaly segmentation and image-level anomaly detection, we report 3 metrics: Area Under Receiver Operator Characteristic curve (AUROC), Average Precision (AP), and $F_{1}$-score at the optimal threshold ($F_{1}$-max). All of these metric are calculated using the scikit-learn library. In addition, we calculate the Intersection over Union (IoU) to more accurately evaluate the anomaly segmentation result.
More training details on anomaly detection methods. In this section, we provide more training details of the comparative anomaly detection methods in Tab .2 and Tab .3 in the main text. For DRAEM [41], GLASS [9], and HVQ-Trans [26], we use the official checkpoints on the MVTec AD dataset, while the others are self-trained due to the lack of official checkpoints. For GLASS, the official foreground masks for the VisA and MVTec 3D AD datasets are not available, so this operation was not used. For PatchCore [32], we use the image size of 256 without center cropping, as some anomalies appear at the edges. For MambaAD [15], we use the provided official checkpoints.
@@
Resource requirement and time consumption. We conduct our training on a NVIDIA Tesla A100 40G GPU sequentially for each product category, which may use about 20G memory. The comparison on time consumption is shown in Tab. 7. For the MVTec AD datasets, our training takes 73 hours, which is shorter than the 249 hours required by AnomalyDiffusion and the 414 hours required by DFMGAN. In terms of inference time, SeaS costs 720 ms per image, which is shorter than the 3830 ms per image required by the Diffusion-based method AnomalyDiffusion. The inference time of the GAN-based method DFMGAN is 48ms per image.
Table 7: Comparison on resource requirement and time consumption.
Methods Training Inference
Overall Time Time (per image)
DFMGAN[11] 414 hours 48 ms
AnomalyDiffusion[17] 249 hours 3830 ms
SeaS 73 hours 720 ms
A.4 More details of the supervised segmentation models
As mentioned in the experiment part, we choose three supervised segmentation models (BiSeNet V2 [40], UPerNet [38], LFD [45]) to verify the validity of the generated image-mask pairs on the downstream supervised anomaly segmentation as well as detection tasks. For BiSeNet V2 and UPerNet, we generally follow the implementation provided by MMsegmentation. For LFD, we also use the official implementation.
In training supervised segmentation models for downstream tasks, we adopt a training strategy of training a unified supervised segmentation model for all classes of products, rather than training separate supervised segmentation models for each class. Experimental results are shown in Tab. 8, which indicate that the performance of the unified supervised segmentation model surpasses that of multiple individual supervised segmentation models.
Table 8: Ablation on the training strategy of supervised segmentation models.
Models Multiple Models Unified Model
AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU
BiSeNet V2 96.00 67.68 65.87 54.11 97.21 69.21 66.37 55.28
UPerNet 96.77 73.88 70.49 60.37 97.87 74.42 70.70 61.24
LFD 93.02 72.97 71.56 55.88 98.09 77.15 72.52 56.47
Average 95.26 71.51 69.31 56.79 97.72 73.59 69.86 57.66
A.5 More ablation studies
Ablation on the Unbalanced Abnormal Text Prompt design
we are able to provide high-fidelity and diverse images for downstream supervised anomaly segmentation tasks, resulting in the best performance in segmentation metrics.
Ablation on the Separation and Sharing Fine-tuning loss
where $A^{l}_{\text{ob}}\in\mathbb{R}^{r\times r\times 1}$ is the cross-attention map corresponding to the normal token <ob>. As shown in Tab. 10, the experimental results demonstrate that, our adopted loss design achieves the best performance in downstream supervised segmentation tasks.
Table 9: Ablation on the Unbalanced Abnormal Text Prompt design.
Prompt AUROC AP $F_{1}$-max IoU
$N^{\prime}$ = 1, $N$ = 1 96.48 63.69 62.50 52.02
$N^{\prime}$ = 1, $N$ = 4 (Ours) 97.21 69.21 66.37 55.28
$N^{\prime}$ = 4, $N$ = 4 96.55 66.28 63.95 54.07
Table 10: Ablation on the Separation and Sharing Fine-tuning loss.
Loss AUROC AP $F_{1}$-max IoU
w/o ST 96.44 67.73 65.23 54.99
with AT 96.42 63.99 62.43 53.36
Ours 97.21 69.21 66.37 55.28
Ablation on the minimum size requirement for training images
In the few-shot setting, for a fair comparison, we follow the common setting in DFMGAN [11] and AnomalyDiffusion [17], i.e., using one-third abnormal image-mask pairs for each anomaly type in training. In this setting, the minimum number of abnormal training images is 2. Once we adopt a 3-shot setting, we need to reorganize the test set. To ensure that the test set is not reorganized for fair comparison, we take 1-shot and 2-shot settings for all anomaly types during training, i.e., $H=1$ and $H=2$, where $H$ is the image number. The results are shown in Tab. 11 and Fig. 9. Observably, the models trained by 1-shot and 2-shot settings still generate anomaly images with decent diversity and authenticity.
Size IS IC-L
$H=1$ 1.790 0.311
$H=2$ 1.794 0.314
$H=\frac{1}{3}\times H_{0}$ 1.876 0.339
Figure 9: Visualization of the ablation study on the minimum size requirement for training images. In the figure, the first row is for generated images, and the second row is for generated masks.
Ablation on the training strategy of SeaS
Table 12: Ablation on training strategy of SeaS.
Strategy IS IC-L
Abnormal-Normal 1.53 0.28
Normal-Abnormal 1.70 0.32
Abnormal&Normal (Ours) 1.88 0.34
In Decoupled Anomaly Alignment (DA) loss, we leverage cross-attention maps from various layers of the U-Net encoder. Specifically, we investigate the impact of integrating different cross-attention maps, denoted as $A^{1}\in\mathbb{R}^{64\times 64}$, $A^{2}\in\mathbb{R}^{32\times 32}$, $A^{3}\in\mathbb{R}^{16\times 16}$ and $A^{4}\in\mathbb{R}^{8\times 8}$. These correspond to the cross-attention maps of the “down-1”, “down-2”, “down-3”, and “down-4” layers of the encoder in U-Net respectively. As shown in Tab. 13, the experimental results demonstrate that, employing a combination of $\{A^{2},A^{3}\}$ for DA loss, achieves the best performance in downstream supervised segmentation tasks.
$A^{l}$ AUROC AP $F_{1}$-max IoU
$l$ = 1, 2, 3 96.42 68.92 66.24 54.52
$l$ = 2, 3, 4 95.71 64.51 62.33 52.46
$l$ = 2, 3 (Ours) 97.21 69.21 66.37 55.28
Ablation on the features for Coarse Feature Extraction
In the coarse feature extraction process, we extract coarse but highly-discriminative features for anomalies from U-Net decoder. Specifically, we investigate the impact of integrating different features, denoted as $F_{1}\in\mathbb{R}^{16\times 16\times 1280}$, $F_{2}\in\mathbb{R}^{32\times 32\times 1280}$, $F_{3}\in\mathbb{R}^{64\times 64\times 640}$ and $F_{4}\in\mathbb{R}^{64\times 64\times 320}$. These correspond to the output feature “up-1”, “up-2”, “up-3”, and “up-4” layers of the encoder in U-Net respectively.
As shown in Fig. 10, we use the output features of the “up-2” and “up-3” layers of the decoder in U-Net, and apply convolution blocks and concatenation operations, then we can obtain the unified coarse feature $\hat{F}\in\mathbb{R}^{64\times 64\times 192}$, which can be used to predict masks corresponding to anomaly images. As shown in Tab. 14, the experimental results demonstrate that, employing a combination of $\{F_{2},F_{3}\}$ for coarse feature extraction, achieves the best performance in the downstream supervised segmentation task.
Figure 10: Visualization of the U-Net decoder features in mask prediction process.
Table 14: Ablation on the features for Coarse Feature Extraction.
$F_{y}$ AUROC AP $F_{1}$-max IoU
$y$ = 1, 2, 3 94.35 63.58 60.54 52.36
$y$ = 2, 3, 4 96.93 67.42 64.26 55.31
$y$ = 2, 3 (Ours) 97.21 69.21 66.37 55.28
Ablation on the features of VAE for Refined Mask Prediction
In the Refined Mask Prediction, we combine the high-resolution features of VAE decoder with discriminative features from U-Net, to generate accurately aligned anomaly image-mask pairs. In addition, we can also use the VAE encoder features as high-resolution features. As shown in Tab. 15, the experimental results show that, using VAE decoder features achieves better performance in downstream supervised segmentation tasks.
Table 15: Ablation on the features of VAE for Refined Mask Prediction.
$F^{res}$ AUROC AP $F_{1}$-max IoU
VAE encoder 96.14 66.26 63.48 54.99
VAE decoder 97.21 69.21 66.37 55.28
Ablation on the normal image supervision for Refined Mask Prediction
In the Refined Mask Prediction branch, we predict masks for normal images as the supervision for the mask prediction. We conduct two sets of experiments: (a) We remove the second and the fourth term in the loss for RMP, i.e., the normal image supervision (short for NIA in Tab. 16); (b) We use the complete form in RMP branch loss, i.e., we use the normal image for supervision, as in Eq. (10):
$\mathcal{L}_{\mathcal{M}}$ $=\mathcal{F}(\hat{M}_{\text{df}},\mathbf{M}_{\text{df}})+\mathcal{F}(\hat{M}_{\text{ob}},\mathbf{M}_{\text{ob}})$ (10)
$\quad+\mathcal{F}(\hat{M}_{\text{df}}^{\prime},\mathbf{M}_{\text{df}}^{\prime})+\mathcal{F}(\hat{M}_{\text{ob}}^{\prime},\mathbf{M}_{\text{ob}}^{\prime})$
As shown in Tab. 16, the experimental results show that, using normal images for supervision achieves better performance in downstream supervised segmentation tasks. We also provide further qualitative results of the effect of normal image supervision (short for NIA in Fig. 11) on MVTec AD.
$F^{res}$ AUROC AP $F_{1}$-max IoU
w/o NIA 96.20 66.03 64.09 53.97
with NIA (Ours) 97.21 69.21 66.37 55.28
Ablation on the Mask Refinement Module
In the Refined Mask Prediction branch, the Mask Refinement Module (MRM) is utilized to generate refined masks. We devise different structures for MRM, as shown in Fig. 12, including Case a): those without conv blocks, Case b): with one conv block, and Case c): with chained conv blocks. As shown in Fig. 13, we find that using the conv blocks in Case b), which consists of two 1 $\times$ 1 convolutions and one 3 $\times$ 3 convolution, helps the model learn the features of the defect area more accurately, rather than focusing on the background area for using one convolution alone in Case a). Based on this observation, we further designed a chained conv blocks structure in Case c), and the acquired features better reflect the defect area. This one-level-by-one level of residual learning helps the model achieve better residual correction results for the defect area features. As shown in Tab. 17 in the Appendix, Case c) improves the performance by + 0.28% on AUROC, + 2.29% on AP and + 2.29% on $F_{1}$-max, + 0.32% on IoU compared with Case b). We substantiate the superiority of the MRM structures that we design, through the results of downstream supervised segmentation experiments.
Figure 13: Visualization of the MRM module intermediate results. The top is for the MRM structure diagram, and the bottom is sequentially for the input image, feature maps of the MRM intermediate process and the predicted mask.
Table 17: Ablation on the Mask Refinement Module.
Model AUROC AP $F_{1}$-max IoU
with MRM (a) 96.75 68.18 64.96 55.51
with MRM (b) 96.93 66.92 64.08 54.96
with MRM (c) 97.21 69.21 66.37 55.28
Ablation on the threshold for mask binarization
In the Refined Mask Prediction branch, we take the threshold $\tau$ for the second channel of refined anomaly masks $\hat{M}_{\text{df}}^{\prime}$ to segment the final anomaly mask. We train supervised segmentation models using anomaly masks with $\tau$ settings ranging from 0.1 to 0.5. As shown in Tab. 18, results indicate that setting $\tau$ = 0.2 yields the best model performance.
Table 18: Ablation on the threshold for mask binarization.
threshold AUROC AP $F_{1}$-max IoU
$\tau$ = 0.1 97.56 65.33 63.38 52.40
$\tau$ = 0.2 (Ours) 97.21 69.21 66.37 55.28
$\tau$ = 0.3 97.20 66.92 64.35 54.68
$\tau$ = 0.4 95.31 63.55 61.97 53.03
$\tau$ = 0.5 94.11 60.85 59.92 50.87
More detailed quantitative results In this section, we report the detailed generation results for each category on the MVTec AD dataset, VisA dataset, and MVTec 3D AD dataset, which are presented in Tab. 19, Tab. 20, and Tab. 21.
Category Crop& Paste [23] SDGAN [28] Defect- GAN [42] DFMGAN [11] Anomaly Diffusion [17] Ours
IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$
bottle 1.43 0.04 1.57 0.06 1.39 0.07 1.62 0.12 1.58 0.19 1.78 0.21
cable 1.74 0.25 1.89 0.19 1.70 0.22 1.96 0.25 2.13 0.41 2.09 0.42
capsule 1.23 0.05 1.49 0.03 1.59 0.04 1.59 0.11 1.59 0.21 1.56 0.26
carpet 1.17 0.11 1.18 0.11 1.24 0.12 1.23 0.13 1.16 0.24 1.13 0.25
grid 2.00 0.12 1.95 0.10 2.01 0.12 1.97 0.13 2.04 0.44 2.43 0.44
hazelnut 1.74 0.21 1.85 0.16 1.87 0.19 1.93 0.24 2.13 0.31 1.87 0.31
leather 1.47 0.14 2.04 0.12 2.12 0.14 2.06 0.17 1.94 0.41 2.03 0.40
metal_nut 1.56 0.15 1.45 0.28 1.47 0.30 1.49 0.32 1.96 0.30 1.64 0.31
pill 1.49 0.11 1.61 0.07 1.61 0.10 1.63 0.16 1.61 0.26 1.62 0.33
screw 1.12 0.16 1.17 0.10 1.19 0.12 1.12 0.14 1.28 0.30 1.52 0.31
tile 1.83 0.20 2.53 0.21 2.35 0.22 2.39 0.22 2.54 0.55 2.60 0.50
toothbrush 1.30 0.08 1.78 0.03 1.85 0.03 1.82 0.18 1.68 0.21 1.96 0.25
transistor 1.39 0.15 1.76 0.13 1.47 0.13 1.64 0.25 1.57 0.34 1.51 0.34
wood 1.95 0.23 2.12 0.25 2.19 0.29 2.12 0.35 2.33 0.37 2.77 0.46
zipper 1.23 0.11 1.25 0.10 1.25 0.10 1.29 0.27 1.39 0.25 1.63 0.30
Average 1.51 0.14 1.71 0.13 1.69 0.15 1.72 0.20 1.80 0.32 1.88 0.34
Table 20: Comparison on IS and IC-LPIPS on VisA. Bold indicates the best performance.
Category DFMGAN [11] AnomalyDiffusion [17] Ours
candle 1.19 0.23 1.28 0.17 1.20 0.12
capsules 1.25 0.22 1.39 0.50 1.58 0.60
cashew 1.25 0.24 1.27 0.26 1.21 0.28
chewinggum 1.33 0.24 1.15 0.19 1.29 0.27
fryum 1.28 0.20 1.20 0.14 1.14 0.21
macaroni1 1.14 0.24 1.15 0.14 1.15 0.18
macaroni2 1.47 0.38 1.56 0.38 1.57 0.39
pcb1 1.12 0.16 1.18 0.35 1.18 0.26
pcb2 1.12 0.26 1.26 0.21 1.25 0.27
pcb3 1.19 0.18 1.21 0.24 1.22 0.21
pcb4 1.21 0.28 1.14 0.25 1.15 0.22
pipe_fryum 1.43 0.32 1.29 0.17 1.31 0.16
Average 1.25 0.25 1.26 0.25 1.27 0.26